OpenStudy (anonymous):

What is (7x − 9) + (3x + 14)?

OpenStudy (anonymous):

10x+5

OpenStudy (anonymous):

Step one. Remove the parenthesis. 7x - 9 + 3x + 14 Step two. Combine like terms. 7x + 3x - 9 + 14 = 10x + 5 The answer is 10x + 5

OpenStudy (anonymous):

\[(7x-9)+(3x+14)=7x-9+3x+14=7x+3x-9+14=10x+5\]

OpenStudy (anonymous):

7x-9+3x+14 7x+3x-9+14 10x+5